OpenStudy (anonymous):

fidn the second derivative of y=4sin^2(3x)

OpenStudy (anonymous):

i know the answer is 144cos^23x -72, but im not sure how to get that

OpenStudy (anonymous):

\[y'=4*2 \sin 3x \cos 3x *3=12*2\sin 3x \cos 3x=12 \sin 6x\] \[y''=12 \cos 6x *6=72 \cos 6x=72\left( 2\cos ^23x-1 \right)=144\cos ^3x-72\]

OpenStudy (anonymous):

correction in the end \[144 \cos ^2 3x-72\]

OpenStudy (anonymous):

thank you!